# Bloch-Beilinson conjectures for Hecke characters and Eisenstein cohomology of Picard surfaces

### v2

@@ 4.4. Automorphic structure of boundary cohomology
defined in Remark 2.7. For any irreducible representation $V_{\lambda}$ of $\mathrm{G}$ of highest weight $\lambda$, denote by $V_{\lambda,K}$ the corresponding local system on $\mathrm{S}_{K}({\mathbb{C}})$. Letting $K$ vary among the compact open subgroups of ${\mathrm{G}}({\mathbb{A}}_{f})$, one obtains a projective system $(\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda,K})_{K}$ of spaces and sheaves, the cohomology of whose projective limit $(\mathrm{S}({\mathbb{C}}),V_{\lambda})$ is such that
@@
as the image of the natural morphism $H_{c}^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})\rightarrow H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})$. Interior cohomology at a finite level is recover by taking $K$-invariants,
@@
Denote by $\partial{\mathrm{S}}_{K}({\mathbb{C}})$ the boundary of the Borel-Serre compactification at any level $K$. From it and from the restriction to $\partial{\mathrm{S}}_{K}({\mathbb{C}})$ of the canonically extended $V_{\lambda}$ we obtain an analogous projective system and, in the projective limit, a ${\mathrm{G}}({\mathbb{A}}_{f})$-module that we denote
$\partial H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})$
@@
$\partial H^{\bullet}(\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda})=\partial H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})^{K}$
@@
[17, Thm. 1] There exists a canonical isomorphism of ${\mathrm{G}}({\mathbb{A}}_{f})$-modules
@@
Let us now specialize the above Theorem to the case of $V_{\lambda}$ being one of the two local systems $V_{k}$ or $V_{k}^{\vee}$ defined in subsection 2.4. If the type of the restriction to $\mathrm{T}^{\prime}$ of a Hecke character $\phi$ of $\mathrm{T}$ is $\chi=(\kappa_{1},\kappa_{2})$, then we denote [^5] The notation is motivated by the fact that in this case, the infinity type of the restriction of $\phi$ to $\mathrm{T}^{\prime}$ is $(\kappa_{2},\kappa_{1}-\kappa_{2})$ (cfr. subsection 3). by
¯$\overline{{\mathbb{Q}}}_{\kappa_{2},\kappa_{1}-\kappa_{2}}$
the $\mathrm{B}({\mathbb{A}}_{f})$-module ¯$\overline{{\mathbb{Q}}}_{\phi}$. The computation of the characters $w\star\lambda$ provided in subsection 2.6 yields then the following.
@@
Let $\lambda=(k,0,k,k)$.
Then there are canonical isomorphisms of ${\mathrm{G}}({\mathbb{A}}_{f})$-modules
$H^{0}(\partial\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\simeq\bigoplus_{\mbox{\emph{\tiny{type}}}(\phi)=\lambda}\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{0,k}$
@@
 The relation between the above Corollary and Corollary 4.3 is the following (a similar relation holds indeed for any system of coefficients $V_{\lambda}$).
@@
under which the spaces on the right hand-side become endowed with a pure $\overline{{\mathbb{Q}}}$-Hodge structure, direct sum of copies of the pure, 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure $I_{\mu,\nu,\overline{{\mathbb{Q}}}}$.
@@ 5. Eisenstein cohomology
In this section, we fix again one of the Picard modular surfaces $\mathrm{S}_{K}$ introduced in Section 2.7 and we describe its Eisenstein cohomology by recalling the results of the article [17] of Harder. This is used at the end of the section in order to construct the first piece of the extensions which are the object of our main theorem, Theorem 1: a sub-Hodge structure $\partial H^{2}(\phi)$ of Eisenstein cohomology, associated to a suitable Hecke character $\phi$.
@@ 5.1. Eisenstein cohomology
as the space of those ${\mathbb{C}}$-valued, $C^{\infty}$-functions $f$ on $\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{A}})$ which satisfy $f(zg)=\omega^{-1}(z)f(g)$ for any $z\in Z(\tilde{\mathrm{G}})({\mathbb{A}})^{0}$, $g\in\mathop{{\rm G}}\nolimits({\mathbb{A}})$. We define the archimedean component
@@
of automorphic forms on $\mathrm{G}$.
@@
which induces an isomorphism of $\mathop{{\rm G}}\nolimits({\mathbb{A}}_{f})$-modules
@@ 5.2. The Eisenstein operator and its constant term
We consider the Hecke character
$\begin{array}[]{ccc}\mathop{|\delta|}\nolimits:\mathrm{T}({\mathbb{A}})&\rightarrow&{\mathbb{R}}^{\times}_{>0}\\ t&\mapsto&|\delta(t)|\,,\end{array}$
where $\delta$ is the element defined in (14). Note that its restriction to $\mathrm{T}^{\prime}$, seen as a Hecke character of $F$, coincides with the idelic norm on ${\mathbb{I}}_{F}$.
One extends this to a map $|\cdot|:\mathrm{G}({\mathbb{A}})\rightarrow{\mathbb{R}}^{\times}$, by considering a suitable maximal compact subgroup $K_{{\mathbb{A}}}$ of $\mathrm{G}({\mathbb{A}})$ and the associated Iwasawa decomposition
@@
and letting ||$\cdot$|| ⋅ | be trivial on $\rm{U}({\mathbb{A}})$ and on $K_{{\mathbb{A}}}$.
Now fix an algebraic Hecke character $\phi$ of $\mathrm{T}$, and an element $\Phi\in I^{*}_{\phi}$. Then it is known ([25]) that for $s\in{\mathbb{C}}$ with large enough real part, the infinite sum [^6] Note that we are not using unitary induction. This explains the absence of the half-sum of positive roots in our formulae.
$\sum\limits_{\gamma\in\mathrm{B}({\mathbb{Q}})\backslash\mathrm{G}({\mathbb{Q}})}\Phi(\gamma g)|\gamma g|^{s}$
@@
Now recall the element $\theta$ of (17). It is well known that the evaluation at $\Psi\in I^{*}_{\phi|\delta|^{s}}$ of the constant term of the operator $\mathop{\rm Eis}\nolimits_{\phi,s}$, formally defined by
$g\mapsto\int\limits_{\rm{U}({\mathbb{Q}})\backslash\rm{U}({\mathbb{A}})}\mathop{\rm Eis}\nolimits_{\phi,s}(\Psi)(ug)\mbox{d}u,\ \ \ g\in\mathrm{G}({\mathbb{A}})$
@@
and that $\mathop{\rm Eis}\nolimits^{*}_{\phi,s}$ has a pole at $s=0$ if and only if this is the case for the above-described constant term. To study whether such a pole occurs, we introduce the following.
@@
    so that $I^{*}_{\phi\mathop{|\delta|}\nolimits^{s}}=\bigotimes\limits_{v}I^{*}_{\phi_{v}\mathop{|\delta_{v}|}\nolimits^{s}}$. Then, the local components $T(\phi_{v},s)$ of the intertwining operator $T(\phi,s)$ are defined, for every $\Psi_{v}\in I^{*}_{\phi_{v}\mathop{|\delta_{v}|}\nolimits^{s}}$ and for every $g_{v}\in\mathop{{\rm G}}\nolimits({\mathbb{Q}}_{v})$, by
@@
The pole at $s=0$ of the intertwining operator is controlled by an object of arithmetic nature. To define it, we use the character $\omega_{F|{\mathbb{Q}}}$ introduced in subsection 2.1, and identify Hecke characters of $\mathrm{T}^{\prime}$ to Hecke characters of $F$ according to the conventions of subsection 3.
@@
    Let $\phi$ be an Hecke character of $\mathrm{T}$, and denote in the same way the Hecke character of $F$ obtained by restricting $\phi$ to $\mathrm{T}^{\prime}$. Denote by $\phi_{{\mathbb{Q}}}$ the restriction of $\phi$ to a Hecke character of ${\mathbb{Q}}$.
@@
    if $v$ corresponds to a finite prime $p$, and as the product of the appropriate $\Gamma$-factors of the occurring $L$-functions, if $v$ is an infinite place. Here $L_{p}(\cdot,\cdot)$ denotes the corresponding factor in the Euler product of the $L$-function under consideration.
@@
Let $v=p$ be an odd prime which does not ramify in $F$ and at which $\phi$ is not ramified. Then consider the spherical function $\Psi^{(0)}_{v}\in I^{*}_{\phi_{v}\mathop{|\delta|}\nolimits_{v}^{s}}$ defined for every $g_{v}=b_{v}k_{v}$ with $b_{v}\in\mathrm{B}({\mathbb{Q}}_{v})$ and $k_{v}\in\mathop{{\rm G}}\nolimits({\mathbb{Z}}_{v})$, by $\Psi^{(0)}_{v}(g_{v})=\phi_{v}\mathop{|\delta|}\nolimits_{v}^{s}(b_{v})\,.$ The statement then follows from Lai’s spelling out of Langlands’ constant term formula in our situation ([23, Sec. 3]), saying that
@@
Hence, we must study the zeroes and poles of the $L$-functions appearing in $c(\phi,s)$. Since the Hecke characters of interest will be those contributing to the description of boundary cohomology of $\mathrm{S}_{K}$, Theorem 4.6 leads us to consider types of the form $w\star\lambda$, where $w$ is an element of the Weyl group of $\mathrm{G}_{F}$ and the characters $w\star\lambda$ have been defined and computed in subsection 2.6. We will only need to consider the case of elements $w$ of length $\geq 2$.
By applying the basic facts about $L$-functions of Hecke characters recalled in subsection 3, we find the following.
@@
  1. Let $\lambda=(k_{1},k_{2})$ be a character of $\mathrm{T}^{\prime}$ and $w$ an element of the Weyl group of $\mathrm{G}_{F}$ of length $\geq 2$. Fix an Hecke character $\phi$ of $F$ of type $w\star\lambda$.
@@
    The $L$-function $L(\phi,s)$ is entire, unless $(k_{1},k_{2})=(0,0)$, $w=$ (1 3) and $\phi(\cdot)=|\cdot|_{{\mathbb{I}}_{F}}^{2}$. In this case, it has a unique first-order pole at $s=-1$.
@@
  1. Let $\lambda=(k_{1},k_{2})$ be a character of $\mathrm{T}^{\prime}$ and $w$ an element of the Weyl group of $\mathrm{G}_{F}$ of length $\geq 2$. Fix an Hecke character $\phi$ of $F$ of type $w\star\lambda$.
@@
    In each of the above two cases, it has a unique first-order pole at $s=-2$.
From the previous two lemmas, we get immediately the following corollary.
@@
[…]

### v4

@@ 4.4. Automorphic structure of boundary cohomology
defined in Remark 2.7. For any irreducible representation $V_{\lambda}$ of $\mathrm{G}$ of highest weight $\lambda$, denoted by $V_{\lambda,K}$ the corresponding local system on $\mathrm{S}_{K}({\mathbb{C}})$. Letting $K$ vary among the compact open subgroups of ${\mathrm{G}}({\mathbb{A}}_{f})$, one obtains a projective system $(\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda,K})_{K}$ of spaces and sheaves, the cohomology of whose projective limit $(\mathrm{S}({\mathbb{C}}),V_{\lambda})$ is such that
@@
as the image of the natural morphism $H_{c}^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})\rightarrow H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})$. Interior cohomology at a finite level is recovered by taking $K$-invariants,
@@
Remark 4.5.
  The $\mathrm{G}({\mathbb{A}}_{f})$-action at infinite level equips the above space of $K$-invariants with a structure of module under the level $K$ Hecke algebra
$\mathcal{C}^{\infty}_{c}(\mathrm{G},K)$
i.e. the algebra of smooth functions with compact support on $\mathrm{G}({\mathbb{A}}_{f})$, bi-invariant under $K$, with product defined by convolution.
Denote by $\partial{\mathrm{S}}_{K}({\mathbb{C}})$ the boundary of the Borel-Serre compactification at any level $K$. From this and from the restriction to $\partial{\mathrm{S}}_{K}({\mathbb{C}})$ of the canonically extended $V_{\lambda}$, we obtain an analogous projective system and, in the projective limit, a ${\mathrm{G}}({\mathbb{A}}_{f})$-module that we denote
$\partial H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})\,.$
@@
$\partial H^{\bullet}(\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda})=\partial H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})^{K}\,.$
@@
[23, Thm. 1] There exists a canonical isomorphism of ${\mathrm{G}}({\mathbb{A}}_{f})$-modules
@@
Let us now specialize the above theorem in the case of $V_{\lambda}$ being one of the two local systems $V_{k}$ or $V_{k}^{\vee}$ defined in Subsection 2.4. If the type of the restriction to $\mathrm{T}^{\prime}$ of a Hecke character $\phi$ of $\mathrm{T}$ is $\chi=(\kappa_{1},\kappa_{2})$, then we denote [^5] The notation is motivated by the fact that, in this case, the infinity type of the restriction of $\phi$ to $\mathrm{T}^{\prime}$ is $(\kappa_{2},\kappa_{1}-\kappa_{2})$ (cfr. Subsection 3). by
$\overline{{\mathbb{Q}}}_{\kappa_{2},\kappa_{1}-\kappa_{2}}$
the $\mathrm{B}({\mathbb{A}}_{f})$-module $\overline{{\mathbb{Q}}}_{\phi}$. The computation of the characters $w\star\lambda$ provided in Subsection 2.6 then yields the following.
@@
Let $\lambda=(k,0,k,k)$. Then there are canonical isomorphisms of ${\mathrm{G}}({\mathbb{A}}_{f})$-modules
$H^{0}(\partial\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\simeq\bigoplus_{\mbox{\emph{\tiny{type}}}(\phi)=id\star\lambda}\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{0,k}$
@@
 The relation between the above Corollary 4.8 and the Corollary 4.3 is the following (a similar relation holds indeed for any system of coefficients $V_{\lambda}$).
@@
under which the spaces on the right hand side become endowed with a pure $\overline{{\mathbb{Q}}}$-Hodge structure, direct sum of copies of the pure, 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure $I_{\mu,\nu,\overline{{\mathbb{Q}}}}$.
@@ 5. Eisenstein cohomology
In this section, we fix again one of the Picard modular surfaces $\mathrm{S}_{K}$ introduced in Section 2.7 and we describe its Eisenstein cohomology by recalling the results of the article [23] of Harder. This is used at the end of the section to construct the first piece of the extensions, which are the object of our Theorem 1. Namely, we construct a sub-Hodge structure $\partial H^{2}(\phi)$ of Eisenstein cohomology, associated to a suitable Hecke character $\phi$.
@@ 5.1. Eisenstein cohomology
as the space of those $\mathbb{C}$}-valued, $C^{\infty}$-functions $f$ on $\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{A}})$ which satisfy
$f(zg)=\omega^{-1}(z)f(g)$
for any $z\in Z(\mathrm{G})({\mathbb{A}})^{0}$, $g\in\mathop{{\rm G}}\nolimits({\mathbb{A}})$. We define the archimedean component
@@
of automorphic forms on $\mathrm{G}$ (see e.g. [19, 6.3]).
@@
(it is the adelic, infinite-level version of [14, VII, Cor. 2.7]). It induces an isomorphism of $\mathop{{\rm G}}\nolimits({\mathbb{A}}_{f})$-modules
@@ 5.2. The Eisenstein operator and its constant term
Consider the element $\delta$ defined in (15), which induces a map, still denoted by $\delta$
$T({\mathbb{A}})\rightarrow{\mathbb{I}}_{F}$
that actually lands in $\mathbb{I}$$_{{\mathbb{Q}}}$. So, we can define the Hecke character
$\begin{array}[]{ccc}\mathop{|\delta|}\nolimits:\mathrm{T}({\mathbb{A}})&\rightarrow&{\mathbb{R}}^{\times}_{>0}\\ t&\mapsto&|\delta(t)|_{{\mathbb{I}}_{{\mathbb{Q}}}},.\end{array}$
Note that its restriction to $\mathrm{T}^{\prime}$, seen as a Hecke character of $F$, coincides with the idelic norm on ${\mathbb{I}}_{F}$.
We extend this to a map $|\cdot|:\mathrm{G}({\mathbb{A}})\rightarrow{\mathbb{R}}^{\times}$, considering a suitable maximal compact subgroup $K_{{\mathbb{A}}}$ of $\mathrm{G}({\mathbb{A}})$ and the associated Iwasawa decomposition
@@
and letting $|\cdot|$ be trivial on $\rm{U}({\mathbb{A}})$ and on $K_{{\mathbb{A}}}$.
Now fix an algebraic Hecke character $\phi$ of $\mathrm{T}$, and an element $\Phi\in I^{*}_{\phi}$. Then it is known ([30]) that for $s\in{\mathbb{C}}$ with a large enough real part, the infinite sum [^6] Note that we are not using unitary induction. This explains the absence of the half-sum of positive roots in our formulae.
[…]
@@
Now recall the element $\theta$ of (18). It is well known that the evaluation of the constant term of the operator $\mathop{\rm Eis}\nolimits_{\phi,s}$ at $\Psi\in I^{*}_{\phi|\delta|^{s}}$, formally defined by
[…]
@@
and that $\mathop{\rm Eis}\nolimits^{*}_{\phi,s}$ has a pole at $s=0$ if and only if this is the case for the constant term described above. To study whether such a pole occurs, we introduce the following.
@@
    so that $I^{*}_{\phi\mathop{|\delta|}\nolimits^{s}}=\bigotimes\limits_{v}I^{*}_{\phi_{v}\mathop{|\delta_{v}|}\nolimits^{s}}$. Then, the local components $T(\phi_{v},s)$ of the intertwining operator $T(\phi,s)$ are defined for all $\Psi_{v}\in I^{*}_{\phi_{v}\mathop{|\delta_{v}|}\nolimits^{s}}$ and for all $g_{v}\in\mathop{{\rm G}}\nolimits({\mathbb{Q}}_{v})$, by
@@
The appearance of a pole at $s=0$ in the intertwining operator is controlled by a quantity of arithmetic nature. To define it, we use the character $\omega_{F|{\mathbb{Q}}}$ introduced in Subsection 2.1, and identify Hecke characters of $\mathrm{T}^{\prime}$ with Hecke characters of $F$ according to the conventions of Subsection 3.1.
@@
    Let $\phi$ be the Hecke character of $\mathrm{T}$, and denote in the same way the Hecke character of $F$ obtained by restricting $\phi$ to $\mathrm{T}^{\prime}$. Denote by $\phi_{{\mathbb{Q}}}$ the restriction of $\phi$ to a Hecke character of $\mathbb{Q}$}.
@@
    if $v$ corresponds to a finite prime $p$, and as the product of the appropriate $\Gamma$-factors of the occurring $L$-functions, if $v$ is an infinite place. Here, $L_{p}(\cdot,\cdot)$ denotes the corresponding factor in the Euler product of the $L$-function under consideration.
@@
Let $v=p$ be an odd prime that does not ramify in $F$ and at which $\phi$ is not ramified. Then consider the spherical function $\Psi^{(0)}_{v}\in I^{*}_{\phi_{v}\mathop{|\delta|}\nolimits_{v}^{s}}$ defined for every $g_{v}=b_{v}k_{v}$ with $b_{v}\in\mathrm{B}({\mathbb{Q}}_{v})$ and $k_{v}\in\mathop{{\rm G}}\nolimits({\mathbb{Z}}_{v})$, by $\Psi^{(0)}_{v}(g_{v})=\phi_{v}\mathop{|\delta|}\nolimits_{v}^{s}(b_{v})\,.$ The statement then follows from Lai’s spelling out of Langlands’ constant term formula in our situation ([28, Sect. 3]), saying that
@@
Hence, we must study the zeros and poles of the $L$-functions appearing in $c(\phi,s)$. Since the Hecke characters of interest will be those contributing to the description of boundary cohomology of $\mathrm{S}_{K}$, Theorem 4.7 leads us to consider types of the form $w\star\lambda$, where $w$ is an element of the Weyl group of $\mathrm{G}_{F}$ and the characters $w\star\lambda$ have been defined and computed in Subsection 2.6. We will only need to consider the case of elements $w$ of length $\geq 2$.
By applying the basic facts about $L$-functions of Hecke characters recalled in Section 3, we find the following.
@@
  1. Let $\lambda=(k_{1},k_{2})$ be a character of $\mathrm{T}^{\prime}$ and $w$ be an element of the Weyl group $W$ of $\mathrm{G}_{F}$ of length $\geq 2$. Fix an Hecke character $\phi$ of $F$ of type $w\star\lambda$.
@@
    The $L$-function $L(\phi,s)$ is entire, unless $(k_{1},k_{2})=(0,0)$, $w=$ (1 3) and $\phi(\cdot)=|\cdot|_{{\mathbb{I}}_{F}}^{2}$. In this case, it has a unique first-order pole at s=1s=-1.
@@
  1. Let $\lambda=(k_{1},k_{2})$ be a character of $\mathrm{T}^{\prime}$ and $w$ be an element of the Weyl group $W$ of $\mathrm{G}_{F}$ of length $\geq 2$. Fix an Hecke character $\phi$ of $F$ of type $w\star\lambda$.
@@
    In each of the above two cases, it has a unique first-order pole at s=2s=-2.
From the previous two lemmas, we immediately get the following corollary.
@@
Let $k$ be a positive integer, put $\lambda=(k,0)$, and let $\phi$ be a Hecke character of $F$ of type $\mbox{\emph{(1\ 2\ 3\ )}}\star\lambda$. Suppose $\phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}}(\cdot)=|\cdot|_{{\mathbb{I}}_{\mathbb{Q}}}^{3}$.